OpenStudy (anonymous):

1.Please help what is the fraction of .666 thanks step by step please

OpenStudy (anonymous):

Quick answer is 2/3

OpenStudy (anonymous):

more complex answer would be to show how as the number of decimal places increases, so more sixes, then the fraction would approach 2/3.

OpenStudy (anonymous):

go On mathway.com it really helps w/ a lot of math questions :)

OpenStudy (anonymous):

\[x=.6666...\\ 10x=6.6666...\\ 9x=6\\ x=\frac{6}{9}=\frac{2}{3}\] is one method

OpenStudy (anonymous):

How did you get the 2/3

OpenStudy (anonymous):

forgot about that method

OpenStudy (anonymous):

i would go with Satelites method

OpenStudy (anonymous):

steps were a) multiply by \(10\) b) subtract \(x\) c) divide by \(9\)

OpenStudy (anonymous):

clear or you want me to do it step by step again?

OpenStudy (anonymous):

Yes please

OpenStudy (deepolisnoob):

Levie wanted an explanation to getting that fraction To convert your decimal to a fraction, you must divide 0.666 by 1, so you'll have \[\frac{0.666}{ 1 }\] You then multiply the numerator and denominator by 10 until you eliminate any decimals... \[\frac{0.666*1000}{1*1000}=\frac{666}{1000}\] You can now simplify numerator and denominator by dividing it by common factors. You can divide both by 2 \[\frac{666/2}{1000/2}=\frac{333}{500}\] That is as far as you can evenly simplify it.

OpenStudy (anonymous):

ok we call \(.6666...\) a variable \(x\) and then solve for it first write \[x=.6666...\]

OpenStudy (anonymous):

then multiply it by \(10\) on the left you get \(10x\)and on the right you move the decimal over one place and get \(6.666...\) so we have \[10x=6.666...\]

OpenStudy (anonymous):

then subtract \(x\) from both sides. \(10x-x=9x\) and \(6.666...-.6666...=6\) as all the repeating decimals go that gives \[9x=6\]

OpenStudy (anonymous):

finally divide by \(9\) and you get \[x=\frac{6}{9}\] which is the same as \[x=\frac{2}{3}\]

OpenStudy (anonymous):

btw \[0.1111...=\frac{1}{9}\\ 0.2222...=\frac{2}{9}\\ 0.3333...=\frac{3}{9}\\ 0.4444...=\frac{4}{9}\] etc

OpenStudy (anonymous):

Thanks all

OpenStudy (anonymous):

yw